OpenStudy (anonymous):

how can i identify the solutions of 3x2 − 13x + 4 = 0

OpenStudy (anonymous):

factor as \[(x-4)(3x-1)=0\] as a start

OpenStudy (anonymous):

then set each factor = 0 and solve you get \[x-4=0\] \[x=4\] or \[3x-1=0\] \[3x=1\] \[x=\frac{1}{3}\]